OpenStudy (anonymous):

Find solutions in the interval [0, 2pi) 2secx+1=cosx

OpenStudy (noorik.):

secx=1/cosx then 2secx-cosx=-1 (2/cosx)-cosx=-1 (2-cos^2(x))/cosx=-1 2-cos^2(x)=-cosx cos^2x-cosx-2=0 (cosx-2)(cosx+1)=0 either cosx=2 (not possible, because -1<=cosx<=1) or cosx=-1, then x= pi

OpenStudy (anonymous):

thank you soooo much.

OpenStudy (noorik.):

u r welcome